OpenStudy (anonymous):

using the powers of 10 notation and conversion factors, calculate the numbers of seconds in the average lifetime 70 years of human

OpenStudy (shane_b):

Where y = years, d = days, h = hours, m = minutes and s = seconds:\[70y * \frac{365d}{1y}*\frac{24h}{1d}*\frac{60m}{1h}*\frac{60s}{1m}=2,207,520,000s=2.20752\space x\space10^{9}s\]